Give an example of a series  ∑an such that  ∑an is not convergent but the sequence (an) converges to 0.


1
Expert's answer
2021-06-15T08:11:50-0400

Consider the harmonic series


"\displaystyle\sum_{i=1}^na_n=\displaystyle\sum_{i=1}^n\dfrac{1}{n}"

The sequence "\{a_n\}" defined by "a_n=\dfrac{1}{n}, n\geq1" converges to "0."

But the series "\displaystyle\sum_{i=1}^n\dfrac{1}{n}" is divergent.